Consider the following aggregate expenditure model of the Canadian economy operating with given wages and other factor prices, price level, interest rates, exchange rates, and expectations:
C = 50 + 0.8YD I = 400 G = 500 T = 0.3Y X = 650 IM = 0.36Y
where C is consumption (the 0.8 term represents the marginal propensity to consume) YD is disposable income, I is investment, G is government spending on goods and services, T is the total value of taxes net of transfers (the 0.3 term represents the net tax rate on national income), X is exports, and IM is imports (the 0.36 term represents the marginal propensity to import).
- Now suppose that the government decides to use its spending power to restore national income to its original level. By how much must the government increase G to restore the original level of national income? What will happen to the government’s budget balance?
1
Expert's answer
2015-02-27T09:04:37-0500
C = 50 + 0.8YD
I = 400
G = 500
T = 0.3Y
X = 650
IM = 0.36Y
(b) Y1 = $1520.9
(c) Y2 = $1420.9
(d) If the government decides to use its spending power to restore national income to its original level, and as G = 500, the government must increase G by 100 to restore the original level of national income.

The new government’s budget balance will be:
BS = 0.3*1520.9 - 600 = -$143.7, so it will decrease by 70, so the budget deficit will increase.